During routine​ screening, a doctor notices that

1717​%

of her adult patients show higher than normal levels of glucose in their

bloodlong dash—a

possible warning signal for diabetes. Hearing​ this, some medical researchers decide to conduct a​ large-scale study, hoping to estimate the proportion to within

55​%

with

9595​%

confidence. How many randomly selected adults must they​ test?
Answer #1

Solution :

Given that,

= 0.17

1 - = 1 - 0.17 = 0.83

margin of error = E = 5% = 0.05

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96 ( Using z table )

Sample size = n = (Z/2 / E)2 * * (1 - )

= (1.96 / 0.05)2 * 0.17 * 0.83

= 216.81

Sample size = 217